# Soliton Amplification in the Korteweg-de Vries Equation by Multiplicative Forcing

### v1

@@ Abstract
We study the stability and dynamics of solitons in the Korteweg de-Vries (KdV) equation with small multiplicative forcing. Forcing breaks the conservative structure of the KdV equation, leading to substantial changes in energy over long time. We show that, for small forcing, the inserted energy is almost fully absorbed by the soliton, resulting in a drastically changed amplitude and velocity. We decompose the solution to the forced equation into a modulated soliton and an infinite dimensional perturbation. Assuming slow exponential decay of the forcing, we show that the perturbation decays at the same exponential rate in a weighted Sobolev norm centered around the soliton.
@@ 1 Introduction
In the absence of forcing ($f(t)\equiv 0$), (1.1) is the well-known KdV equation; a well-studied dispersive PDE that first appeared in the description of shallow water waves in a longitudinal canal [13]. Among its most notable features is the existence of soliton solutions $u(t,x)=\phi_{c}(x-ct)$ of the form
@@
With $\mathcal{N}$, $\mathcal{H}$ and other KdV-invariants undergoing slow (but eventually large) changes, we may expect significant consequences for the propagation of the solitons (1.2). Using (1.4) and the relation $\mathcal{N}[\phi_{c}]=6c^{3/2}$, we can heuristically predict that the amplitude of a soliton starting at $c(0)>0$ will approximately evolve according to
@@
$u(0,x)=\phi_{c_{*}}(x)+\overline{v}_{*}(x)$
for some $c_{*}>0$, where $\overline{v}_{*}\in H^{2}$ and $e^{wx}\overline{v}_{*}\in H^{1}$ are suitably small for some weight $w\in(0,\sqrt{c_{*}}/3)$. If the forcing term $f$ is assumed to be exponentially decaying, then the solution actually converges to a limiting wave profile in $H^{1}$ with an exponential weight centered around the soliton. Our main interest in pursuing this line of work is to open up rigorous stability results for solitons undergoing large amplitude changes. As such, we view this work as a step towards understanding the stability of solitons under more general perturbations, such as stochastic forcing [18], as well as the stability of KdV-like quasi-solitons such as micropterons and nanopterons in systems/lattices with periodic structure [4, 14, 5, 11, 6]. Indeed, these gradually decrease in amplitude over time when perturbed, due to the presence of small oscillatory tails that interact with the perturbation.
@@
Pick $c_{*},E_{\mathrm{max}}>0$, $w\in(0,\sqrt{c_{*}}/3)$ and $p\in[0,\tfrac{1}{4})$. There exist a weight $w_{\infty}\in(0,w)$ and constants $\delta_{1},C_{1},c_{\mathrm{min}},c_{\mathrm{max}}>0$ with $c_{\mathrm{min}}<c_{*}<c_{\mathrm{max}}$ such that the following holds true. For each $E\in(0,E_{\mathrm{max}}]$, each $\epsilon\in(0,1]$ that satisfies
@@
there exist modulation functions $c,\xi\in C^{1}({\mathbb{R}}^{+};{\mathbb{R}})$ that satisfy the following properties:
@@
    $\sup_{t\geq 0}\big{|}c(t)-c_{\mathrm{ap}}(t)\big{|}+\sup_{t\geq 0}\left|\xi(t)-\xi_{\mathrm{ap}}(t)\right|\leq$ $2C_{1}E\epsilon^{1-4p}+3C_{1}\|\overline{v}_{*}\|_{H^{1}_{w}}+2C_{1}(E+\delta_{1})\|\overline{v}_{*}\|_{H^{1}}.$
@@
    The integrability of $f$ ensures that $c(t)$ and $c^{-1}(t)$ remain bounded, which prevents technical complications. Firstly, because it allows us to construct bounds that do not depend on $c(t)$. Secondly, it guarantees that we can apply a minimal exponential weight on the modulated soliton $\phi_{c(t)}$.
@@
    The assumption $\overline{v}_{*}\in H^{2}\cap H^{1}_{w}$ fits the well-posedness results established in [17, Appendix A]. It would be interesting to see if this assumption can be relaxed to $\overline{v}_{*}\in L^{2}$ by following the arguments of [16].
@@
A related result is available in [20] for a different forcing type, where the authors use a factorisation technique that is not available in the current setting.
@@
The paper is organized as follows. In Section 2, we derive a system of modulation equations that governs the behavior of the soliton amplitude $c(t)$, the position $\xi(t)$ and remainder $v(t)$. Then, in Section 3, we introduce the function spaces that are central to our stability argument, and assert stability and smoothing properties of the operator $\mathcal{L}_{c}$ on these spaces. We proceed by establishing control of the remainder $v$ over short-timescales by adapting a Duhamel argument of [17], and introduce the notion of time-varying weights. Thereafter, in Section 5, we show that the remainder $v$ can be controlled over long time-scales in a weighted norm. The evolution of unweighted norms of $v$ is then analysed in Section 6. We finally provide the proof of Theorem 1.1 in Section 7.
Acknowledgements
Both authors acknowledge support from the Dutch Research Council (NWO) (grant 613.009.137).
@@ 2 Modulation system
In order to track how constants depend on the system parameters and the various choices that we make, we collect the various assumptions that we make throughout the paper in a number of labelled ‘settings’. The first of these relates to the global parameters $(c_{*},E_{\mathrm{max}},w)$ and the initial condition for (2.1).
@@
This result is established in [17, Appendix A] for $f(t)\equiv 0$ by modifying a well-posedness result of Kato [12]. To see that the arguments for local well-posedness (i.e. small $T>0$) remain valid upon including the forcing term $\epsilon f(\gamma t)$, we note that (2.1) is equivalent to a time-dependent KdV equation. Indeed, $u(t)$ solves (2.1) if and only if $w(t)=e^{-(\epsilon/\gamma)\int_{0}^{\gamma t}f(s){\mathrm{d}}s}u(t)$ solves
$w_{t}=-\partial_{x}^{3}w-2e^{(\epsilon/\gamma)\int_{0}^{\gamma t}f(s){\mathrm{d}}s}w\partial_{x}w.$
The arguments for global well-posedness (i.e. arbitrary $T>0$) rely on an a-priori bound for the $H^{2}$-norm. We show here that such a bound remains available upon including the forcing term $\epsilon f(\gamma t)$. Indeed, writing $u(t)=u$ for some $t\geq 0$ and using (2.1) we compute that
@@
and
@@
[…]
which leads to (1.4) and the a-priori bound
$\mathcal{H}[u(t)]\leq\mathcal{H}[u(0)]e^{3\epsilon/\gamma\int_{0}^{\gamma t}f(s){\mathrm{d}}s}.$
Introducing the higher-order quantity
$\mathcal{E}_{2}[u]=\int_{\mathbb{R}}(\partial_{x}^{2}u)^{2}-\tfrac{10}{3}u(\partial_{x}u)^{2}{\mathrm{d}}x,$
@@
$\partial_{t}\mathcal{E}_{2}[u]=$ $2\epsilon f(\gamma t)\mathcal{E}_{2}[u]+\tfrac{20}{3}\langle u,u_{x}^{3}\rangle+\epsilon f(\gamma t)\int_{{\mathbb{R}}}\tfrac{10}{3}u(\partial_{x}u)^{2}+\tfrac{20}{3}u^{2}\partial_{x}^{2}u\ {\mathrm{d}}x,$
through which one arrives at an a-priori bound on $\mathcal{E}_{2}[u(t)]$ and hence $\|u(t)\|_{H^{2}}$. ∎
@@
and $\xi,c$ are time-dependent modulation parameters. The existence, uniqueness and continuous time-dependence of this decomposition is guaranteed by Lemma A.2 as long as $\|\overline{v}(t)\|_{L^{2}_{w}}$ is kept below some constant $\delta_{2}>0$. Based on this decomposition, we introduce a phase-shift parameter $\Omega$ through
@@
Item 1 follows from (2.4) by substituting $y=\alpha(t)x=\sqrt{c_{0}/c(t)}$. In the same way, item 2 follows from (2.5). Finally, we compute the norms
@@
which is necessary to ensure that $\langle v,\phi_{c_{0}}\rangle=\langle v,\zeta_{c_{0}}\rangle=0$ and equivalently $\langle\overline{v},\phi_{c(t)}\rangle=\langle\overline{v},\zeta_{c(t)}\rangle=0$. The matrix $K(v)$ is invertible in case $\|v\|_{L^{2}_{w}}$ is suitably small, since $K(0)$ is invertible and $v\mapsto\det K(v)$ is continuous from $L^{2}_{w}$ to ${\mathbb{R}}$. Consequently, the system (2.7), (2.9) is well-defined as long as $\|v(t)\|_{L^{2}_{w}}$ remains suitably bounded.
@@
In particular, the term $x\partial_{x}[\phi_{c_{0}}+v]$ in (2.8) lies in $L_{b}^{2}$, since there exists a constant $C>0$ for which we have
@@ 3 Linear stability on weighted spaces
Here, $\eta_{c}^{1}$ and $\eta_{c}^{2}$ are linear combinations of $\phi_{c}$ and $\zeta_{c}$ that are defined in (A.2). As such, the subspace of $L^{2}_{w}$ characterised by (1.10) corresponds to $\ker P_{c}\subseteq L^{2}_{w}$. We review this classical result in Appendix A, where it is stated as Theorem A.1.
@@
for $b_{-}<w<b_{+}$, some constant $C>0$ and functions $g$ for which the above quantity is well-defined. This leads us to introduce the notion of asymmetrically-weighted spaces. For every $\mathbf{w}=(w_{-},w_{+})\in{\mathbb{R}}^{2}$, we introduce the weighted space
@@
The following result asserts that the stability and smoothing properties of $\{e^{\mathcal{L}_{c}t}\}_{t\geq 0}$ provided in Theorem A.1 extend to the asymmetrically-weighted spaces.
@@
there exists a constant $M>0$ such that for all $g\in L_{\mathbf{w}}^{2}$, $t>0$ and $k\in\{0,1\}$ we have
@@
If $\mathbf{w},\mathbf{b}\in{\mathbb{R}}^{2}$ satisfy $b_{-}<w_{-}$, $b_{+}>w_{+}$ and $g\in L^{2}_{\mathbf{b}}$, then $g,xg\in L^{2}_{\mathbf{w}}$ and
@@ 4 Short-time control
In this section, we establish control over the perturbation in the original-frame in asymmetrically-weighted spaces over short time-scales. Using our rescaled frame description (2.7) on short time-scales leads to problematic complications arising from the term $x\partial_{x}$ in (2.8). Instead, we rely on classical results valid for small amplitude fluctuations. We follow the argument of Pego & Weinstein [17, Proposition 6.1], which uses the evolution
@@
    Given $T>0$, the function $\overline{v}\in C([0,T],H^{2})\cap C^{1}([0,T],H^{-1})$ satisfies (4.1) while the function $\alpha\in C^{1}([0,T],{\mathbb{R}})$ solves (2.9). In addition, we have the inclusion
@@
Our main result in this section provides short-time control over $\overline{v}$. More precisely, on time-intervals where $\overline{v}$ and the fluctuations of $c$ are small enough, the growth of $\overline{v}$ measured in the $H^{1}_{\overline{\mathbf{w}}}$-norm can be explicitly controlled by the small forcing amplitude $\epsilon>0$ and the length of the time-interval $\delta>0$.
@@
Assumption C2 implies that $t\to\tfrac{w_{+}(t)}{{\alpha}(t)}$ is decreasing and $t\to\tfrac{w_{-}(t)}{{\alpha}(t)}$ is increasing on $[0,T]$, which essentially means that the weight-functions retain their monotonicity after rescaling. Since $w_{-}$ and $w_{+}$ are evaluated at $t+s/2$ in (4.5), it furthermore follows that there is a lower bound on their absolute growth rate, in the sense that
@@
Let us write $v=v(t)$ for brevity. Setting out to control the matrix inverse $K^{-1}(v)$ in (2.9), we note that
@@
$\|\zeta_{c_{0}}\|^{2}_{L^{2}_{-\mathbf{b}}}=$ $\int_{-\infty}^{0}e^{-2b_{-}x}\zeta_{c_{0}}^{2}(x){\mathrm{d}}x+\int_{0}^{\infty}e^{-2b_{+}x}\zeta_{c_{0}}^{2}(x){\mathrm{d}}x$
$\leq$ $\int_{-\infty}^{0}\zeta_{c_{0}}^{2}(x){\mathrm{d}}x+\frac{\|\zeta_{c_{0}}\|^{2}_{L^{\infty}}}{2b_{+}}$
and we thus have
@@
for all $(i,j)\in\{1,2\}^{2}$ and some constant $\tilde{C}_{3}>0$. In case $\delta_{6}\leq b_{+}^{1/2}\tfrac{\tilde{C}_{1}}{\tilde{C}_{3}}$, it follows via (4.7) that $\|K^{-1}(v)\|_{\mathrm{op}}\leq\tilde{C}_{2}$. We proceed by estimating
@@
$|\langle\phi_{c_{0}}+{v},\phi_{c_{0}}\rangle|+|\langle\phi_{c_{0}}+{v},\zeta_{c_{0}}\rangle|\leq$ $6c_{0}^{3/2}+9+\|\phi_{c_{0}}\|_{L^{2}_{-\mathbf{b}}}\|v\|_{L^{2}_{\mathbf{b}}}+\tilde{C}_{3}b_{+}^{-1/2}\|v\|_{L^{2}_{\mathbf{b}}}.$
@@
$|\alpha_{t}|+|\Omega_{t}|\leq$ $\alpha_{\mathrm{max}}\epsilon|f(\gamma t)|\tilde{C}_{2}\Big{(}6c_{0}^{3/2}+9+\|\phi_{c_{0}}\|_{L^{2}_{-\mathbf{b}}}\|v\|_{L^{2}_{\mathbf{b}}}+\tilde{C}_{3}b_{+}^{-1/2}\|v\|_{L^{2}_{\mathbf{b}}}\Big{)}$
$+\alpha_{\mathrm{min}}^{-2}\tilde{C}_{2}\tilde{C}_{4}\|v\|^{2}_{L^{2}_{\mathbf{b}}}.$
@@
We write $\overline{v}=\overline{v}(t_{\diamond}+s)$ and estimate the various components in (4.2). Firstly, we have

### v2

@@ Abstract
We study the stability and dynamics of solitons in the Korteweg-de Vries (KdV) equation with small multiplicative forcing. Forcing breaks the conservative structure of the KdV equation, leading to substantial changes in energy over long times. We show that, for small forcing, the inserted energy is almost fully absorbed by the soliton, resulting in a drastically changed amplitude and velocity. We decompose the solution to the forced equation into a modulated soliton and an infinite dimensional perturbation. Assuming slow exponential decay of the forcing, we show that the perturbation decays at the same exponential rate in a weighted Sobolev norm centered around the soliton.
@@ 1 Introduction
In the absence of forcing ($f(t)\equiv 0$), (1.1) is the well-known KdV equation: a well-studied dispersive PDE that first appeared in the description of shallow water waves in a longitudinal canal [14]. Among its most notable features is the existence of soliton solutions $u(t,x)=\phi_{c}(x-ct)$ of the form
@@
With $\mathcal{N}$, $\mathcal{H}$, and other KdV-invariants undergoing slow (but eventually large) changes, we may expect significant consequences for the propagation of the solitons (1.2). Using (1.4) and the relation $\mathcal{N}[\phi_{c}]=6c^{3/2}$, we can heuristically predict that the amplitude of a soliton starting at $c(0)>0$ will approximately evolve according to
@@
$u(0,x)=\phi_{c_{*}}(x)+\overline{v}_{*}(x)$ (1.7)
for some $c_{*}>0$, where $\overline{v}_{*}\in H^{2}$ and $e^{wx}\overline{v}_{*}\in H^{1}$ are suitably small for some weight $w\in(0,\sqrt{c_{*}}/3)$. If the forcing term $f$ is assumed to be exponentially decaying, then the solution actually converges to a limiting wave profile in $H^{1}$ with an exponential weight centered around the soliton. Our main interest in pursuing this line of work is to open up rigorous stability results for solitons undergoing large amplitude changes. As such, we view this work as a step towards understanding the stability of solitons under more general perturbations, such as stochastic forcing [19], as well as the stability of KdV-like quasi-solitons such as micropterons and nanopterons in systems/lattices with periodic structure [4, 15, 5, 11, 6]. Indeed, these gradually decrease in amplitude over time when perturbed due to the presence of small oscillatory tails that interact with the perturbation.
@@
Pick $c_{*},E_{\mathrm{max}}>0$, $w\in(0,\sqrt{c_{*}}/3)$, and $p\in[0,\tfrac{1}{4})$. There exist a weight $w_{\infty}\in(0,w)$ and constants $\delta_{1},C_{1},c_{\mathrm{min}},c_{\mathrm{max}}>0$ with $c_{\mathrm{min}}<c_{*}<c_{\mathrm{max}}$ such that the following holds true. For each $E\in(0,E_{\mathrm{max}}]$, each $\epsilon\in(0,1]$ that satisfies
@@
there exist modulation functions $c,\xi\in C^{1}({\mathbb{R}}^{+};{\mathbb{R}})$ associated to the solution $u$ of (1.1) with (1.7) that satisfy the following properties:
@@
    $\sup_{t\geq 0}\big{|}c(t)-c_{\mathrm{ap}}(t)\big{|}+\sup_{t\geq 0}\left|\xi(t)-\xi_{\mathrm{ap}}(t)\right|\leq$ $2C_{1}E\epsilon^{1-4p}+3C_{1}\|\overline{v}_{*}\|_{H^{1}_{w}}$
    $+2C_{1}(E+\delta_{1})\|\overline{v}_{*}\|_{H^{1}}.$
@@
    The integrability of $f$ ensures that $c(t)$ and $c^{-1}(t)$ remain bounded, which prevents technical complications. First, it allows us to construct bounds that do not depend on $c(t)$. Second, it guarantees that we can apply a minimal exponential weight on the modulated soliton $\phi_{c(t)}$.
@@
    The assumption $\overline{v}_{*}\in H^{2}\cap H^{1}_{w}$ fits the well-posedness results established in [18, Appendix A]. It would be interesting to see if this assumption can be relaxed to $\overline{v}_{*}\in L^{2}$ by following the arguments of [17].
@@
Two related results for different forcing types are available in [12, 21]. The result in [12] deals with finite time stability, and in [21], the authors use a factorization technique that is not available in the current setting.
@@
The paper is organized as follows. In Section 2, we derive a system of modulation equations that governs the behavior of the soliton amplitude $c(t)$, the position $\xi(t)$, and remainder $v(t)$. Then, in Section 3 we introduce the function spaces that are central to our stability argument, and assert stability and smoothing properties of the operator $\mathcal{L}_{c}$ on these spaces. We proceed by establishing control of the remainder $v$ over short-timescales by adapting a Duhamel argument of [18], and introduce the notion of time-varying weights. Thereafter, in Section 5 we show that the remainder $v$ can be controlled over long time-scales in a weighted norm. The evolution of unweighted norms of $v$ is then analyzed in Section 6. We finally provide the proof of Theorem 1.1 in Section 7.
@@ 2 Modulation system
In order to track how constants depend on the system parameters and the various choices that we make, we collect the various assumptions that we make throughout the paper in a number of labeled ‘settings’. The first of these relates to the global parameters $(c_{*},E_{\mathrm{max}},w)$ and the initial condition for (2.1).
@@
This result is established in [18, Appendix A] for $f(t)\equiv 0$ by modifying a well-posedness result of Kato [13]. To see that the arguments for local well-posedness (i.e. small $T>0$) remain valid upon including the forcing term $\epsilon f(\gamma t)$, we note that (2.1) is equivalent to a time-dependent KdV equation. Indeed, $u(t)$ solves (2.1) if and only if $z(t)=e^{-(\epsilon/\gamma)\int_{0}^{\gamma t}f(s){\mathrm{d}}s}u(t)$ solves
$z_{t}=-\partial_{x}^{3}z-2e^{(\epsilon/\gamma)\int_{0}^{\gamma t}f(s){\mathrm{d}}s}z\partial_{x}z.$
The arguments for global well-posedness (i.e. arbitrary $T>0$) rely on an a priori bound for the $H^{2}$-norm. We show here that such a bound remains available upon including the forcing term $\epsilon f(\gamma t)$. Indeed, writing $u(t)=u$ for some $t\geq 0$ and using (2.1), we compute that
@@
leading to the identity
$\|u(t)\|^{2}_{L^{2}}=e^{2(\epsilon/\gamma)\int_{0}^{\gamma t}f(s){\mathrm{d}}s}\|u(0)\|^{2}_{L^{2}}.$
Moving on to the first derivative, an application of the Gagliardo-Nirenberg inequality yields
$\|\partial_{x}u\|_{L^{2}}^{2}=2\mathcal{H}[u]+\tfrac{2}{3}\int_{{\mathbb{R}}}u^{3}{\mathrm{d}}x\leq 2\Big{|}\mathcal{H}[u]\Big{|}+C\|u\|_{L^{2}}^{5/2}\|\partial_{x}u\|_{L^{2}}^{1/2}.$
This inequality is of the form $x^{4}\leq A+B|x|$ with $A,B\geq 0$, from which we may conclude $x^{4}\leq 2(A+B^{4/3})$, i.e.
$\|\partial_{x}u\|_{L^{2}}^{2}\leq 4\Big{|}\mathcal{H}[u]\Big{|}+2C^{4/3}\|u\|_{L^{2}}^{10/3}.$ (2.4)
We then compute
@@
=\ = $\epsilon f(\gamma t)(3\mathcal{H}[u]-\tfrac{1}{2}\|\partial_{x}u\|_{L^{2}}^{2}),$
which leads to
$\Big{|}\partial_{t}\mathcal{H}[u]\Big{|}\leq 5\epsilon|f(\gamma t)|\Big{|}\mathcal{H}[u]\Big{|}+\epsilon|f(\gamma t)|C^{4/3}\|u\|_{L^{2}}^{10/3}$
and, via an application of Grönwall’s inequality, to the a priori bound
$\Big{|}\mathcal{H}[u(t)]\Big{|}\leq\Big{(}\mathcal{H}[u(0)]+\frac{\epsilon}{\gamma}C^{4/3}\int_{0}^{\gamma t}|f(s)|{\mathrm{d}}s\sup_{0\leq s\leq t}\|u(s)\|_{L^{2}}^{10/3}\Big{)}e^{5\epsilon/\gamma\int_{0}^{\gamma t}|f(s)|{\mathrm{d}}s}.$
Via (2.4), this provides an a priori bound on $\|\partial_{x}u(t)\|_{L^{2}}^{2}$. Using the fact that the integral
$\mathcal{E}_{2}[u]=\int_{\mathbb{R}}(\partial_{x}^{2}u)^{2}-\tfrac{10}{3}u(\partial_{x}u)^{2}+\tfrac{5}{9}u^{4}{\mathrm{d}}x$
is conserved for the unperturbed KdV flow, we may use the bound $\|u\|_{L^{4}}^{4}\leq 4\|u\|_{H^{1}}^{4}$ to estimate
$\|\partial_{x}^{2}u\|_{L^{2}}^{2}=\mathcal{E}_{2}[u]+\tfrac{10}{3}\int_{{\mathbb{R}}}u(\partial_{x}u)^{2}{\mathrm{d}}x-\tfrac{5}{9}\int_{{\mathbb{R}}}u^{4}{\mathrm{d}}x\leq\Big{|}\mathcal{E}_{2}[u]\Big{|}+\tfrac{10}{3}\|u\|_{H^{1}}^{3}+\tfrac{20}{9}\|u\|_{H^{1}}^{4}.$
@@
$\partial_{t}\mathcal{E}_{2}[u]=$ $2\epsilon f(\gamma t)\mathcal{E}_{2}[u]+\epsilon f(\gamma t)\int_{{\mathbb{R}}}-\tfrac{10}{3}u(\partial_{x}u)^{2}+\tfrac{10}{9}u^{4}\ {\mathrm{d}}x,$
so that
$\Big{|}\partial_{t}\mathcal{E}_{2}[u]\Big{|}\leq$ $2\epsilon|f(\gamma t)|\Big{|}\mathcal{E}_{2}[u]\Big{|}+\epsilon|f(\gamma t)|\Big{(}\tfrac{10}{3}\|u\|_{H^{1}}^{3}+\tfrac{10}{9}\|u\|_{H^{1}}^{4}\Big{)},$
through which one arrives at an a priori bound on $\mathcal{E}_{2}[u(t)]$, and hence $\|u(t)\|_{H^{2}}$. ∎
@@
and $\xi,c$ are time-dependent modulation parameters. The existence, uniqueness, and continuous time-dependence of this decomposition is guaranteed by Lemma A.2 as long as $\|\overline{v}(t)\|_{L^{2}_{w}}$ is kept below some constant $\delta_{2}>0$. Based on this decomposition, we introduce a phase-shift parameter $\Omega$ through
@@
Item 1 follows from (2.5) by substituting $y=\alpha(t)x=\sqrt{c_{0}/c(t)}x$. In the same way, item 2 follows from (2.6). Finally, we compute the norms
@@
which is necessary to ensure that $\langle v,\phi_{c_{0}}\rangle=\langle v,\zeta_{c_{0}}\rangle=0$, and equivalently $\langle\overline{v},\phi_{c(t)}\rangle=\langle\overline{v},\zeta_{c(t)}\rangle=0$. The matrix $K(v)$ is invertible in case $\|v\|_{L^{2}_{w}}$ is suitably small, since $K(0)$ is invertible and $v\mapsto\det K(v)$ is continuous from $L^{2}_{w}$ to ${\mathbb{R}}$. Consequently, the system (2.8), (2.10) is well-defined as long as $\|v(t)\|_{L^{2}_{w}}$ remains suitably bounded.
@@
In particular, the term $x\partial_{x}[\phi_{c_{0}}+v]$ in (2.9) lies in $L_{b}^{2}$ since there exists a constant $C>0$, for which we have
@@ 3 Linear stability on weighted spaces
Here, $\eta_{c}^{1}$ and $\eta_{c}^{2}$ are linear combinations of $\phi_{c}$ and $\zeta_{c}$ that are defined in (A.2). As such, the subspace of $L^{2}_{w}$ characterized by (1.11) corresponds to $\ker P_{c}\subseteq L^{2}_{w}$. We review this classical result in Appendix A, where it is stated as Theorem A.1.
@@
for $b_{-}<w<b_{+}$, some constant $C>0$, and functions $g$ for which the above quantity is well-defined. This leads us to introduce the notion of asymmetrically-weighted spaces. For every $\mathbf{w}=(w_{-},w_{+})\in{\mathbb{R}}^{2}$, we introduce the weighted space
@@
The following result asserts that the stability and smoothing properties of $\{e^{\mathcal{L}_{c}t}\}_{t\geq 0}$ provided in Theorem A.1 extend to asymmetrically-weighted spaces.
@@
there exists a constant $M>0$ such that for all $g\in L_{\mathbf{w}}^{2}$, $t>0$, and $k\in\{0,1\}$, we have
@@
If $\mathbf{w},\mathbf{b}\!\!\in\!{\mathbb{R}}^{2}$ satisfy $b_{-}\!<\!w_{-}$, $b_{+}\!>\!w_{+}$ and $g\!\in\!L^{2}_{\mathbf{b}}$, then $g,xg\!\in\!L^{2}_{\mathbf{w}}$, and
@@ 4 Short-time control
In this section, we establish control over the perturbation in the original frame in asymmetrically-weighted spaces over short time-scales. Using our rescaled frame description (2.8) on short time-scales leads to problematic complications arising from the term $x\partial_{x}$ in (2.9). Instead, we rely on classical results valid for small amplitude fluctuations. We follow the argument of Pego & Weinstein [18, Proposition 6.1], which uses the evolution
@@
    Given $T>0$, the function $\overline{v}\in C([0,T],H^{2})\cap C^{1}([0,T],H^{-1})$ satisfies (4.1), while the function $\alpha\in C^{1}([0,T],{\mathbb{R}})$ solves (2.10). In addition, we have the inclusion
@@
Our main result in this section provides short-time control over $\overline{v}$. More precisely, on time intervals where $\overline{v}$ and the fluctuations of $c$ are small enough, the growth of $\overline{v}$ measured in the $H^{1}_{\overline{\mathbf{w}}}$-norm can be explicitly controlled by the small forcing amplitude $\epsilon>0$ and the length of the time interval $\delta>0$.
@@
Assumption C2 implies that $t\to\tfrac{w_{+}(t)}{{\alpha}(t)}$ is decreasing and $t\to\tfrac{w_{-}(t)}{{\alpha}(t)}$ is increasing on $[0,T]$, which essentially means that the weight-functions retain their monotonicity after rescaling. Since $w_{-}$ and $w_{+}$ are evaluated at $t+s/2$ in (4.5), it furthermore follows that there is a lower bound on their absolute growth rate in the sense that
@@
$|\alpha_{t}(t)+\tfrac{2}{3}\alpha(t)\epsilon f(\gamma t)|+|\Omega_{t}(t)-\tfrac{2}{3}c_{0}^{-1/2}\alpha(t)\epsilon f(\gamma t)|$
$\leq C_{6}\epsilon|f(\gamma t)|b_{+}^{-1/2}\|v(t)\|_{L^{2}_{\mathbf{b}}}+C_{6}\|v(t)\|^{2}_{L^{2}_{\mathbf{b}}},$ (4.6)
and hence
@@
Let us write $v=v(t)$ for brevity, and rewrite (2.10) as
$\begin{bmatrix}{\alpha}_{t}+\tfrac{2}{3}\alpha(t)\epsilon f(\gamma t)\\ {\Omega}_{t}-\tfrac{2}{3}c_{0}^{-1/2}\alpha(t)\epsilon f(\gamma t)\end{bmatrix}=$ $-{\alpha}\epsilon{f}(\gamma t)K^{-1}({v})\begin{bmatrix}\langle{v},\phi_{c_{0}}\rangle\\ \langle{v},\zeta_{c_{0}}\rangle\end{bmatrix}$
$-\alpha f(\gamma t)(K^{-1}({v})-K^{-1}(0))\begin{bmatrix}\langle\phi_{c_{0}},\phi_{c_{0}}\rangle\\ \langle\phi_{c_{0}},\zeta_{c_{0}}\rangle\end{bmatrix}$
$-{\alpha}^{-2}K^{-1}({v})\begin{bmatrix}\langle N({v}),\phi_{c_{0}}\rangle\\ \langle N({v}),\zeta_{c_{0}}\rangle\end{bmatrix}.$ (4.8)
Here we have used that
$K^{-1}(0)\begin{bmatrix}\langle\phi_{c_{0}},\phi_{c_{0}}\rangle\\ \langle\phi_{c_{0}},\zeta_{c_{0}}\rangle\end{bmatrix}=\begin{bmatrix}\frac{2}{3}\\ -\frac{2}{3}c_{0}^{-1/2}\end{bmatrix}.$
Setting out to control $K^{-1}(v)$, we note that
@@
$\|\zeta_{c_{0}}\|^{2}_{L^{2}_{-\mathbf{b}}}\!=\!\int_{-\infty}^{0}e^{-2b_{-}x}\zeta_{c_{0}}^{2}(x){\mathrm{d}}x\!+\!\int_{0}^{\infty}e^{-2b_{+}x}\zeta_{c_{0}}^{2}(x){\mathrm{d}}x\leq\int_{-\infty}^{0}\zeta_{c_{0}}^{2}(x){\mathrm{d}}x+\frac{\|\zeta_{c_{0}}\|^{2}_{L^{\infty}}}{2b_{+}}$
and have
@@
for all $(i,j)\in\{1,2\}^{2}$ and some constant $\tilde{C}_{3}>0$. in case $\delta_{6}\leq b_{+}^{1/2}\tfrac{\tilde{C}_{1}}{\tilde{C}_{3}}$, it follows via (4.9) that $\|K^{-1}(v)\|_{\mathrm{op}}\leq\tilde{C}_{2}$. Turning to the term $K^{-1}({v})-K^{-1}(0)$ in (4.8), we note that
$\|K^{-1}({v})-K^{-1}(0)\|_{\mathrm{op}}$ $=\|K^{-1}(0)\big{(}K(v)-K(0)\big{)}K^{-1}(v)\|_{\mathrm{op}}$
$\leq\|K^{-1}(0)\|_{\mathrm{op}}\|K(v)-K(0)\|_{\mathrm{op}}\|K^{-1}(v)\|_{\mathrm{op}}$
$\leq\tilde{C}_{4}b_{+}^{-1/2}\|v\|_{L^{2}_{\mathbf{b}}}.$
We proceed by estimating
@@
$|\langle{v},\phi_{c_{0}}\rangle|+|\langle{v},\zeta_{c_{0}}\rangle|\leq$ $\|\phi_{c_{0}}\|_{L^{2}_{-\mathbf{b}}}\|v\|_{L^{2}_{\mathbf{b}}}+\tilde{C}_{3}b_{+}^{-1/2}\|v\|_{L^{2}_{\mathbf{b}}}.$
@@
$|\alpha_{t}+\tfrac{2}{3}\alpha(t)\epsilon f(\gamma t)|+|\Omega_{t}-\tfrac{2}{3}c_{0}^{-1/2}\alpha(t)\epsilon f(\gamma t)|$
$\leq\alpha_{\mathrm{max}}\epsilon|f(\gamma t)|\tilde{C}_{2}\Big{(}\|\phi_{c_{0}}\|_{L^{2}_{-\mathbf{b}}}\|v\|_{L^{2}_{\mathbf{b}}}+\tilde{C}_{3}b_{+}^{-1/2}\|v\|_{L^{2}_{\mathbf{b}}}\Big{)}$
$\quad+\alpha_{\mathrm{max}}\epsilon|f(\gamma t)|\tilde{C}_{4}\Big{(}\|\phi_{c_{0}}\|_{L^{2}}^{2}+|\langle\phi_{c_{0}},\zeta_{c_{0}}\rangle|\Big{)}b_{+}^{-1/2}\|v\|_{L^{2}_{\mathbf{b}}}+\alpha_{\mathrm{min}}^{-2}\tilde{C}_{2}\tilde{C}_{5}\|v\|^{2}_{L^{2}_{\mathbf{b}}}.\qed$
@@
We write $\overline{v}=\overline{v}(t_{\diamond}+s)$ and estimate the various components in (4.2). First, we have